HETAV PATEL

Grade 11,

A horse is drinking water from a cubical container of side 1m kept on the ground .The height of stomach of horse from the ground level is 2m .How much work will be done by horse to drink the entire water of cubical block

A horse is drinking water from a cubical container of side 1m kept on the ground .The height of stomach of horse from the ground level is 2m .How much work will be done by horse to drink the entire water of cubical block

Grade:11

1 Answers

Subhankar Biswas
4 years ago
W=mgh
So,
m=d×V=1000×(1)^3=1000kg
So,W=1000×10×2 J
W=20000J=20kJ.
Therefore the total Work done by the horse in drinking the water is 20kJ.